# Cost-of-capital valuation with risky assets

### v1

@@ 6 Conclusions
The results of this paper may be generalized in various directions. For instance, it could be interesting to investigate the sensitivity of the resulting quantities with respect to model uncertainty, both on the dynamics of the financial market and the distribution of insurance claims. Furthermore, although the independence assumption between insurance and financial risk is a common assumption in practice for many non-life business lines, it can be interesting to see how introducing dependence may compromise this trade-off of introducing additional risk for the prospect of higher investment returns. Also, to extend the analysis to a multi-period framework including explicit runoffs in long-tailed business would be of interest (see e.g. [25, 22], possibly also under consideration of a changing realized cost-of-capital rate, cf. [1]). Such extensions are left for future research.
@@ Appendix A Proofs
$\frac{\partial}{\partial y}F_{X_{1}-rZ_{1}}(y)=\int_{y+r(1+w)}^{\infty}f_{wS_{1}}\bigg(\frac{x-y}{r}+w-1\bigg)\frac{1}{r}dF_{X_{1}}(x)>0.$
@@
$\frac{\partial}{\partial r}F_{X_{1}-rZ_{1}}(y)=\int_{y+r(1+w)}^{\infty}f_{wS_{1}}\bigg(\frac{x-y}{r}+w-1\bigg)\frac{x-y}{r^{2}}dF_{X_{1}}(x)>0.$
Since $r\mapsto g(r):=F_{X_{1}-rZ_{1}}(0)$ is continuous and strictly increasing, $g(0)=F_{X_{1}}(0)<1-\alpha$ and $g(r)\downarrow 1$ as $r\to\infty$, there exists a unique $r>0$ solving $F_{X_{1}-rZ_{1}}(0)=1-\alpha$.
@@
Suppose that $\widetilde{X}_{1}\geq_{\text{st}}X_{1}$. Let $\widetilde{R}_{0}$ be the unique solution to $\rho(\widetilde{R}_{0}Z_{1}-\widetilde{X}_{1})=0$. Suppose that $\widetilde{R}_{0}<R_{0}$. We will see that this assumption leads to a contradiction. By the monotonicity property of $\operatorname{VaR}_{\alpha}$ and $\operatorname{ES}_{\alpha}$ and the independence of $X_{1}$ and $\widetilde{X}_{1}$ of $Z_{1}$,
@@
Suppose now that $\widetilde{Z}_{1}\geq_{\text{st}}Z_{1}$. Let $\widetilde{R}_{0}$ be the unique solution to $\rho(\widetilde{R}_{0}\widetilde{Z}_{1}-X_{1})=0$. Suppose that $\widetilde{R}_{0}>R_{0}$. We will see that this assumption leads to a contradiction. By the monotonicity property of $\operatorname{VaR}_{\alpha}$ and $\operatorname{ES}_{\alpha}$ and the independence of $X_{1}$ and $\widetilde{X}_{1}$ of $Z_{1}$,
@@
Since (16) and (17) are contradictory we conclude that the assumption $\widetilde{R}_{0}>R_{0}$ is false. Hence, $\widetilde{R}_{0}\leq R_{0}$ which completes the proof of the second statement. ∎
Proof of Proposition 2.3.
(i): The proof reuses some arguments from the proof of Proposition 1 in [10]. By Proposition 2.2, $\widetilde{X}_{1}\geq_{\text{st}}X_{1}$ implies $\widetilde{R}_{0}\geq R_{0}$. Note that
@@
(ii) For $\widetilde{Z}_{1}\geq_{\text{st}}Z_{1}$ one has $\widetilde{R}_{0}\leq R_{0}$ (cf. Proposition 2.2). Then
@@
On the other hand, for $\mu>\sigma\Phi^{-1}(1-\alpha)$, there exists an $x>0$ with $f_{\rho}(x)=0$. From (7) it follows that
$(\mu^{2}-\sigma^{2}\Phi^{-1}(1-\alpha)^{2})\,R_{0}^{2}+(-2\mu\gamma)\,R_{0}+(\gamma^{2}-\nu^{2}\Phi^{-1}(1-\alpha)^{2})=0,$
which has two real solutions (ensured by the condition $\mu>\sigma\Phi^{-1}(1-\alpha)$). If $\gamma\leq\nu\Phi^{-1}(1-\alpha)$, then the quadratic equation has a unique positive solution $R_{0}>0$. If $\gamma>\nu\Phi^{-1}(1-\alpha)$, then the quadratic equation has two positive solutions $R_{0}>0$. However, only $R_{0}$ given by (8) solves $f_{\rho}(x)=0$. ∎
@@
Proof of Proposition 4.1.
Note that $\operatorname{VaR}_{\alpha}(R_{0}Z_{1}-X_{1})=0$ is equivalent to
$1-\alpha=\mathbb{P}(X_{1}-R_{0}Z_{1}\leq 0)=\mathbb{P}(X_{1}/Z_{1}\leq R_{0}).$
Since $X_{1}/Z_{1}\sim\text{LN}(m_{x}-m_{z},s_{z}^{2}+s_{x}^{2})$ the expression for $R_{0}$ follows. The expression for $V_{0}$ follows from (2) and (3). ∎

### v2

@@ 5 Conclusions
The results of this paper may be generalized in various directions. For instance, it could be interesting to investigate the sensitivity of the resulting quantities with respect to model uncertainty, both on the dynamics of the financial market and the distribution of insurance claims. Furthermore, although the independence assumption between insurance and financial risk is not an uncommon assumption in practice for certain non-life business lines, in a future study we intend to investigate to what extent dependence may compromise this trade-off of introducing additional risk for the prospect of higher investment returns. Also, to extend the analysis to a multi-period framework including explicit runoffs in long-tailed business would be of interest (see e.g. [26, 23], possibly also under consideration of a changing realized cost-of-capital rate, cf. [1]). Such extensions are left for future research.
@@ Appendix A Proofs
$\frac{\partial}{\partial y}F_{X_{1}-rZ_{1}}(y)=\int_{y+r(w-1)}^{\infty}f_{wS_{1}}\bigg(\frac{x-y}{r}+w-1\bigg)\frac{1}{r}dF_{X_{1}}(x)>0.$
@@
$\frac{\partial}{\partial r}F_{X_{1}-rZ_{1}}(y)=\int_{y+r(w-1)}^{\infty}f_{wS_{1}}\bigg(\frac{x-y}{r}+w-1\bigg)\frac{x-y}{r^{2}}dF_{X_{1}}(x)>0.$
Since $r\mapsto g(r):=F_{X_{1}-rZ_{1}}(0)$ is continuous and strictly increasing, $g(0)=F_{X_{1}}(0)<1-\alpha$ and $g(r)\uparrow 1$ as $r\to\infty$, there exists a unique $r>0$ solving $F_{X_{1}-rZ_{1}}(0)=1-\alpha$.
@@
(i): Suppose that $\widetilde{X}_{1}\geq_{\text{st}}X_{1}$. Let $\widetilde{R}_{0}$ be the unique solution to $\rho(\widetilde{R}_{0}Z_{1}-\widetilde{X}_{1})=0$. Suppose that $\widetilde{R}_{0}<R_{0}$. We will see that this assumption leads to a contradiction. By the monotonicity property of $\operatorname{VaR}_{\alpha}$ and $\operatorname{ES}_{\alpha}$ and the independence of $X_{1}$ and $\widetilde{X}_{1}$ of $Z_{1}$,
@@
(ii): Suppose now that $\widetilde{Z}_{1}\geq_{\text{st}}Z_{1}$. Let $\widetilde{R}_{0}$ be the unique solution to $\rho(\widetilde{R}_{0}\widetilde{Z}_{1}-X_{1})=0$. Suppose that $\widetilde{R}_{0}>R_{0}$. We will see that this assumption leads to a contradiction. By the monotonicity property of $\operatorname{VaR}_{\alpha}$ and $\operatorname{ES}_{\alpha}$ and the independence of $X_{1}$ and $\widetilde{X}_{1}$ of $Z_{1}$,
@@
Since (15) and (16) are contradictory we conclude that the assumption $\widetilde{R}_{0}>R_{0}$ is false. Hence, $\widetilde{R}_{0}\leq R_{0}$ which completes the proof of the second statement.
(iii): The proof reuses some arguments from the proof of Proposition 1 in [10]. By Statement (i), $\widetilde{X}_{1}\geq_{\text{st}}X_{1}$ implies $\widetilde{R}_{0}\geq R_{0}$. Note that
@@
(iv) By Statement (ii), $\widetilde{Z}_{1}\geq_{\text{st}}Z_{1}$ implies $\widetilde{R}_{0}\leq R_{0}$. Then
@@
Write $r(w):=R_{0}^{w}$. For all $w$, $\mathbb{P}(X\leq r(w)(wS_{1}+1-w))=1-\alpha$. Suppose that $X_{1}$ has a density function $f_{X_{1}}$. From
$\int F_{X_{1}}(r(w)(ws+1-w))dF_{S_{1}}(s)=1-\alpha\text{ for all }w$
follows that the derivative of the left-hand side with respect to $w$ is zero. The assumptions allow differentiating under the integral sign, which gives, for all $w$,
$\int f_{X_{1}}(r(w)(ws+1-w))\Big(r^{\prime}(w)(ws+1-w)+r(w)(s-1)\Big)dF_{S_{1}}(s)=0.$
Let
$h(w,s):=f_{X_{1}}(r(w)(ws+1-w))\Big(r^{\prime}(w)(ws+1-w)+r(w)(s-1)\Big)$
and note that the assumptions imply that $|h(w,s)|\leq c_{1}s+c_{2}$ for constants $c_{1},c_{2}>0$. Applying the Dominated Convergence Theorem gives
$0$ $=\lim_{w\to 0}\int f_{X_{1}}(r(w)(ws+1-w))\Big(r^{\prime}(w)(ws+1-w)+r(w)(s-1)\Big)dF_{S_{1}}(s)$
$=\int\lim_{w\to 0}f_{X_{1}}(r(w)(ws+1-w))\Big(r^{\prime}(w)(ws+1-w)+r(w)(s-1)\Big)dF_{S_{1}}(s)$
$=f_{X_{1}}(r(0))\Big(r^{\prime}(0)+r(0)(\mathbb{E}[S_{1}]-1)\Big)$
Hence, $r^{\prime}(0)=-r(0)(\mathbb{E}[S_{1}]-1)<0$.
Set $\psi(w,s):=r(w)(ws+1-w)$.
$\frac{d}{dw}\mathbb{E}[(r(w)(wS_{1}+1-w)-X_{1})^{+}]$
$\quad=\frac{d}{dw}\int_{0}^{\infty}\int_{0}^{\psi(w,s)}\big(\psi(w,s)-x\big)f_{X_{1}}(x)dxdF_{S_{1}}(s)$
$\quad=\int_{0}^{\infty}\int_{0}^{\psi(w,s)}\frac{\partial}{\partial w}\psi(w,s)f_{X_{1}}(x)dxdF_{S_{1}}(s)$
$\quad=\int_{0}^{\infty}F_{X_{1}}\big(\psi(w,s)\big)\Big(r^{\prime}(w)(ws+1-w)+r(w)(s-1)\Big)dF_{S_{1}}(s)$
Since $\lim_{w\to 0}\psi(w,s)=r(0)$, applying the Dominated Convergence Theorem gives
$\lim_{w\to 0}\frac{d}{dw}\mathbb{E}[(r(w)(wS_{1}+1-w)-X_{1})^{+}]$
$\quad=\int_{0}^{\infty}F_{X_{1}}\big(r(0)\big)\Big(r^{\prime}(0)+r(0)(s-1)\Big)dF_{S_{1}}(s)$
$\quad=F_{X_{1}}(r(0))(r^{\prime}(0)+r(0)(\mathbb{E}[S_{1}]-1))$
$\quad=0.$
Since $V_{0}^{w}=R_{0}^{w}-C_{0}^{w}$ the proof is complete. ∎
Proof of Proposition 3.1.
@@
On the other hand, for $\mu>\sigma\Phi^{-1}(1-\alpha)$, there exists an $x>0$ with $f_{\rho}(x)=0$. From (7) it then follows that the function
$g(x)=(\mu^{2}-\sigma^{2}\Phi^{-1}(1-\alpha)^{2})\,x^{2}+(-2\mu\gamma)\,x+(\gamma^{2}-\nu^{2}\Phi^{-1}(1-\alpha)^{2})$
has two real zeroes. If $\gamma\leq\nu\Phi^{-1}(1-\alpha)$, then $g(x)$ has a unique positive zero $R_{0}>0$. If $\gamma>\nu\Phi^{-1}(1-\alpha)$, the function $g(x)$ has two positive zeroes that are candidates for $R_{0}$. However, $f_{\rho}(R_{0})=0$ is equivalent to $\Phi^{-1}(1-\alpha)\sqrt{\sigma^{2}R_{0}^{2}+\nu^{2}}=\mu R_{0}-\gamma$, so $R_{0}$ must satisfy $\mu R_{0}-\gamma\geq 0$. Since $g(0)>0$ and $g(\gamma/\mu)<0$, only $R_{0}$ given in (8) fulfills $\mu R_{0}-\gamma\geq 0$. ∎